OpenStudy (unklerhaukus):

\[u=\int\frac{\text dx}{x^2\sqrt{1+x^2}}\]

OpenStudy (unklerhaukus):

\[\text{let }x=\tan\theta\]\[\text dx=\sec^2\text d\theta\] \[u=\int\frac{\sec^2\theta\cdot \text d\theta}{\tan^2\theta\sqrt{1+\tan^2\theta}}\]\[u=\int\frac{\sec^2\theta\cdot \text d\theta}{\tan^2\theta\sec\theta}\]\[u=\int\frac{\sec\theta\cdot \text d\theta}{\tan^2\theta}\]

OpenStudy (unklerhaukus):

im stuck

OpenStudy (anonymous):

perhaps change everything into cos and sin?

OpenStudy (anonymous):

\[ \frac{\sec (\theta )}{\tan ^2(\theta )}=\cot (\theta ) \csc (\theta ) \] The answe is - csx(\theta) + C

OpenStudy (anonymous):

\[ -\csc(\theta) \]

OpenStudy (anonymous):

+ Constant

OpenStudy (unklerhaukus):

\[u=\int\frac{\sec\theta\cdot \text d\theta}{\tan^2\theta}\]\[=\int\frac 1{\cos \theta}\times\frac{\cos^2\theta}{\sin^2\theta}\text d\theta\]\[=\int\frac{\cos\theta}{\sin^2\theta}\text d\theta\]\[=\int\cot\theta\csc\theta\text d\theta\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

then use the integration table

OpenStudy (anonymous):

i hope i never get an integral like this on the test... i've never seen this rule before

OpenStudy (unklerhaukus):

how do i integrate \[=\int\cot\theta\csc\theta\text d\theta\] integration by parts?

OpenStudy (anonymous):

not entirely sure

OpenStudy (anonymous):

i just see it in my book on the front page with all the basic integrations

OpenStudy (unklerhaukus):

how can i show that @eliassaab

OpenStudy (anonymous):

Didn't you learn how to find the derivative of csc(x)? That is easy.

OpenStudy (anonymous):

Derivative of csc(x) is - csc(x) cot(x)

OpenStudy (unklerhaukus):

i cant remember how

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{1}{sin(x)}\]= \[sin(x)^{-1}\]=-1(sin(x)^{-2}*cos(x)=-1/sin(x)*cos(x)/sin(x)

OpenStudy (anonymous):

=-csc(x)cot(x)

OpenStudy (unklerhaukus):

so i have to use integration by parts , twice

OpenStudy (anonymous):

now let cot(x) = u u=csc(x) du=-csc(x)cot(x) \[\int{}{}-du\]

OpenStudy (anonymous):

\[-\int{}{}du=-x+c=-csc(x)+c\]

OpenStudy (anonymous):

i find that super strange lol but it works

OpenStudy (anonymous):

i've never sen a problem like that amazingly

OpenStudy (anonymous):

don't read cot(x)=u .... i meant csc(x)

OpenStudy (unklerhaukus):

it just a tiny part in a larger question

OpenStudy (anonymous):

are you a teacher or just have homework that you have to do on the computer?!?

OpenStudy (unklerhaukus):

this is just question from my DE subject textbook form second year university, i am opting to work in LaTeX because it is neater

OpenStudy (unklerhaukus):

im not a teacher just a physics student

OpenStudy (anonymous):

lol you use a lot of function that my teacher doesn't use... such as mew (t) for your integrating factor

OpenStudy (anonymous):

that's interesting though... dose your teacher give you not a lot of questions... it seems like that question is a doozie

OpenStudy (unklerhaukus):

actually my book uses \(R(x)\) for integrating factors , i have stated using \(\mu(x)\) because it seams to be what most people use , the other common symbol used for the integrating factor i have seen is \(IF(x)\)

OpenStudy (unklerhaukus):

most of the questions im by book have ended up taking over a page of latex,

OpenStudy (unklerhaukus):

i guess i can use this standard integral \[\int\sin^m(x)\cos^n(x)\text dx=\frac{\sin^{m+1}(x)\cos^{n-1}}{m+n}+\frac{n-1}{m+n}\int\sin(x)^m\cos^{n-2}(x)\text dx\]

OpenStudy (unklerhaukus):

m=-2 n=1

OpenStudy (unklerhaukus):

\[=\int\frac{\cos\theta}{\sin^2\theta}\text d\theta=\]\[\int\sin^{-2}(\theta)\cos(\theta)\text d\theta=\frac{\sin^{-1}(\theta)}{-1}=-\csc(\theta)+C\]

OpenStudy (unklerhaukus):

how can i show these are equal\[u=-\csc(\theta)+C\]\[u=-\frac{\sqrt{x^2+1}}{x}\]

OpenStudy (anonymous):

why not let u= 1+x^2, try that

OpenStudy (unklerhaukus):

oh i forgot \[x=\tan\theta\] \[\theta=\arctan x\]

OpenStudy (anonymous):

figure it out?

OpenStudy (unklerhaukus):

\[=-\csc(\arctan x)\]

OpenStudy (unklerhaukus):

how can i simplify the trig functions

OpenStudy (anonymous):

you let tan(theta)=x correct... well if you construct your triangle for you trig sub

OpenStudy (anonymous):